OpenStudy (anonymous):

If the circumference of a circle is 12, and the angle measure of an arc is 60o, which is the length of the arc?

hartnn (hartnn):

From the circumference, can you first find the radius ?

hartnn (hartnn):

use the formula \(Circumference ~ C =2 \pi r \)

hartnn (hartnn):

C = 12 find r

OpenStudy (anonymous):

im confused

hartnn (hartnn):

12 = 2\(\pi\)r r = 12\(\div (2\pi)\) = ...?